Andrew is an analyst at a wealth management firm. One of his clients holds a $10,000 portfolio that consists of four stocks.Suppose, based on the earnings consensus of stock analysts, Andrew expects a return of 12.14% from the portfolio with the newdrop down 1 options: 0.26%, 0.20%, 0.32%, or 0.30%

drop down 2 options: decrease or increase

drop down 3 options: increase or decrease
Answer #1

Please do Upvote if you are served. Feel free to reach out in the comments

Cheers!!!

Answer:

We know from CAPM model (Re) = Rf+ Beta * Market Risk premium

Given Rf= 6% and Market risk premium = 7.5%,

Return of Atteric = 6+0.6*7.5 = 10.5%

Return of Arthur = 6+1.5*7.5 = 17.25%

Reuturn of Lobster = 6 + 1.3*7.5 = 15.75%

Return of Baque = 6+0.5*7.5 = 9.75%

After all shares of atteric are replaced by Baque, weight of Baque = 30+35 =65%=0.65

=> New return of portfolio = 0.2 * 17.25 + 0.15*15.75 + 0.65 * 9.75 = 12.15%

Change = 12.41 - 12.15 = 0.26%

The stock expert expects 12.14% which is equal to the calculated return of 12,15%, we can say that the stocks are Fairly valued.

If the Atteric is replaced with company X stock, then total portfolio beta would be higher and the required return from the portflio would be higher